Anna has a pocket of dimes and quarters if she has 10 coins worth 1.45 how many of her coins are quarters

Answer by harpazo(655)

Multiply 1.45 by 100 to remove the decimal value.
So, 1.45 x 10 = 145.

d + q = 10
10d + 25q = 145
===============
d = 10 - q
10(10- q) + 25q = 145
100 - 10q + 25q = 145
100 + 15q = 145
15q = 145 - 100
15q = 45
q = 45/15
q = 3
The number of quarters she has is 3.